OpenStudy (anonymous):

The cost to rent a construction crane is $750 per day plus $250 per hour of use. What is the maximum number of hours the crane can be used each day if the rental cost is not to exceed $2500 per day?

OpenStudy (anonymous):

PLEASE SHOW WORK :]

OpenStudy (anonymous):

C= 750 +250h 2500=750 + 250h 1750 = 250h h= 7 7 hours a day

OpenStudy (anonymous):

you need help? answer me in my question posted the names dlopez21413

OpenStudy (anonymous):

Alright then I'm happy with my response.